OpenStudy (anonymous):

ln 1/sqrt(e) = x, how do I calculate x?

OpenStudy (anonymous):

ln (1/sqrt(e)) = x ?

OpenStudy (anonymous):

you know that ln(a/b) = lna - lnb ?

OpenStudy (anonymous):

\[\ln \frac{ 1 }{ \sqrt{e} } = x\]

OpenStudy (anonymous):

can you do it now ?

OpenStudy (anonymous):

Damn, thanks lol. I'm blind.

OpenStudy (anonymous):

or you can convert 1/sqrt(e) to e^-0.5

OpenStudy (anonymous):

it will turn out to be the same answer ofcurse

Parth (parthkohli):

\[\ln\left|{1 \over \sqrt{e}}\right| \implies \ln1 -\ln\sqrt{e} \implies 0 - {1 \over 2}\]

OpenStudy (anonymous):

Thanks =))

OpenStudy (anonymous):

:)